# Stellar cores live long and prosper in cuspy dark matter halos

### v1

@@ Abstract
The existence of cuspy or cored centers of dark matter halos is a crucial discriminant between different dark matter models. It has recently been claimed based on dynamical arguments that perfectly cored stellar systems cannot survive inside cuspy dark matter halos, which would make the observation of stellar cores in ultra-faint dwarf galaxies, where dark matter cores cannot form through baryonic processes, a direct falsification of the cold dark matter paradigm. Here, we use idealized simulations to show explicitly that cored stellar systems like those observed in dwarf galaxies can be stable within cuspy dark matter halos over at least several Hubble times. We also demonstrate that observations of ultra-faint dwarf galaxies cannot distinguish mildly positive, flat, or negative inner density slopes, further precluding the dynamical inference of the gravitational potential from the stellar configuration.
@@
software: gadget-4 (springel_simulating_2021), pynbody (pynbody), NumPy (harris_array_2020), SciPy (virtanen_scipy_2020), Matplotlib (hunter_computing_2007), Stan (standevelopmentteam2018), CmdStanPy (standevelopmentteam2018), Arviz (kumar2019).
@@
1 Introduction
Dwarf galaxies, and in particular, the dwarf-spheroidal (dSph) and ultra-faint dwarf (UFD) galaxies surrounding the Milky Way, are among the Universe’s most dark-matter-dominated objects. Both in their abundance and in their internal structure, they are key to understanding the nature of dark matter (DM).
In the standard, cold dark matter (CDM) model, collisionless DM forms so-called “cuspy” halos, characterized by a divergent inner density slope that is well described by a Navarro–Frenk–White (NFW) density profile (Navarro-1997). Observational tests of this prediction on scales of individual galaxies have yielded mixed results. While rotation curves derived for disk galaxies tend to favor “cored” halos of near-uniform density in their central regions (e.g., oh11; li20), the diversity of observed rotation curves includes individual cases that favor NFW-like cusps (e.g., oman15).
@@ 1 Introduction
Various techniques have been used to infer the mass distributions within “classical” dSphs (stellar masses in the range $\sim 10^{4}\mbox{--}10^{7}\,\mathrm{M}_{\sun}$) from photometric, spectroscopic, and in some cases proper motion data — again with mixed results. Several studies conclude that the most luminous dSphs (e.g., Fornax and Sculptor) harbor large DM cores, based on Jeans models (Read_2019; Hayashi_2020), orbit-based Schwarzschild models (Jardel_2012), direct modeling of the phase-space distribution function (DF) (Pascale_2018), and statistical analysis of chemodynamically distinct subpopulations (Battaglia_2008; Walker_2011; Amorisco_2011; Agnello_2014). Others, analyzing the same spectroscopic data sets, claim insufficient constraining power to distinguish DM cores from CDM cusps (e.g., Strigari_2017; Strigari_2018; Zhu_2016). Including Hubble Space Telescope proper motion data that resolve the internal velocity dispersions of the Draco dSph, Vitral_2025 constrain any DM core to have radius $\lesssim 1\,\mathrm{kpc}$. Several studies applying homogeneous analyses to all available dSphs conclude that the most luminous ones tend to favor cores, while the least luminous ones favor cusps (e.g., Read_2019; Hayashi_2020; Hayashi_2023).
@@
This scenario is consistent with results of some cosmological and hydrodynamical simulations, which show that supernova feedback from sufficiently vigorous star formation can alter the structure of CDM halos, inducing the formation of cores in what would originally have been cuspy halos (Navarro-1996; mashchenko2008; Pontzen-2012) hosting the most luminous dSphs. This process of astrophysical core formation within CDM halos is expected to become inefficient on scales of the smallest and least luminous UFD galaxies, where feedback from the formation of their paltry stellar populations would have little effect on the structure of the host DM halo (e.g., madau2014; dicintio2014; onorbe2015; Orkney-2021).
@@
Thus, the detection of cores in UFDs would directly rule out collisionless CDM. However, the fact that UFDs contain so few stars, usually in a single population, makes the analysis that is already difficult in their brighter counterparts even more so (see, e.g., simon2019, for a review).
@@
In a series of papers, almeida_EIM_2024 proposed a new approach to determine the DM density profile of UFDs, requiring only their (projected) surface brightness profiles and applying the Eddington inversion method (almeida2023) to infer the gravitational potential. They concluded that an isotropic stellar population cannot have a cored density profile within a cuspy DM halo, because it implies that the phase-space DF, derived using the Eddington inversion method, has negative values. In particular, almeida_stellar_2024 claim that the combined surface brightness profiles of six UFDs require a cored density profile and, thus, rule out a cuspy halo, and Almeida-2025 explore these in the context of self-interacting DM.
@@
Here, we test this proposition in two parts. Firstly, we explicitly test the stability of cored stellar systems inside cuspy DM halos, using idealized numerical simulations. We find that they can be stable in cuspy halos, provided that the two are initially set up in equilibrium. Secondly, we test the ability to constrain the stellar density profile of UFDs using a finite number of stars. We find that, in particular, the inner slope is degenerate, even before considering observational limitations.
2 Numerical methods
@@
Traditional cosmological simulations lack both the mass and the spatial resolution to simultaneously model the DM halo and resolve the stellar systems of UFDs. Furthermore, they require gravitational softening, which could artificially induce or maintain cores in the stellar components (e.g., Power-2003).
@@
2.1 Initial conditions
We generate the initial conditions (ICs) used to model the dwarf galaxies using the DF method following hilz2012. Our fiducial system consists of a stellar component of mass $M_{\star}=5\times 10^{3}\,\mathrm{M}_{\sun}$ embedded in a DM halo with a virial mass of $M_{\mathrm{vir}}=5\times 10^{8}\,\mathrm{M}_{\sun}$. The DM component is modeled by an NFW density profile:
@@ 2.1 Initial conditions
$\rho_{\mathrm{S,NFW}}=\frac{M_{\mathrm{vir}}}{4\pi r_{\mathrm{S,NFW}}^{3}}\\ \times\left[\log\left(1+\frac{r_{\mathrm{vir}}}{r_{\mathrm{S,NFW}}}\right)-\frac{r_{\mathrm{vir}}}{r_{\mathrm{S,NFW}}+r_{\mathrm{vir}}}\right]^{-1},$ (2)
@@
The stellar component is modeled by a Plummer density profile (binney2008):
@@
The DF $f_{i}$ is computed for both components from their respective density profiles $\rho_{i}$ using Eddington’s formula (binney2008):
@@
where $\sigma_{r}$, $\sigma_{\theta}$, and $\sigma_{\phi}$ are the radial, polar, and azimuthal velocity dispersions, respectively, is zero throughout.
2.2 Simulations
We run idealized $N$-body simulations using the gadget-4 code (springel_simulating_2021), compiled using the Fast Multipole Method with two order multipoles, hierarchical time integration, and an external gravitational field created by a static NFW potential. The stellar distribution is set up to be in equilibrium with the NFW halo at $t=0$. In practice, we add to the equations of motion in gadget-4 an additional acceleration term due to an external NFW profile of mass $M_{\mathrm{ext}}=M_{\mathrm{vir}}$:
$\bm{a}_{\mathrm{NFW}}=\frac{G}{r^{3}}M_{\mathrm{ext}}\\ \times\left[\log\left(1+\frac{r}{r_{\mathrm{S,NFW}}}\right)-\frac{r}{r_{\mathrm{S,NFW}}}\left(1+\frac{r}{r_{\mathrm{S,NFW}}}\right)^{-1}\right]\\ \times\left(\log(1+c)-\frac{c}{1+c}\right)\bm{x},$ (6)
@@ 2.2 Simulations
We create and perform 10 different realizations of our fiducial system. The number of realizations is chosen to mimic the number of UFDs used in the analysis of almeida_stellar_2024. To estimate the effect of the external potential, we also deviate its virial mass, $M_{\mathrm{ext}}$, (and radius) from that of one of the initial fiducial systems, $M_{\mathrm{IC}}$, and perform seven simulations with an unequal mass ratio $M_{\mathrm{ext}}/M_{\mathrm{IC}}$. In other words, in these simulations, the initial stellar system is explicitly out of equilibrium with the potential in which it will be evolved.
We investigate the effect of softening by performing simulations with softening lengths of 1, 10 and 100 pc. There was no clear distinction between these values, and we chose a softening of 10 pc to be used for this work.
To study both the short-time evolution and the long-term stability, the simulations are analyzed at $t=0\,\mathrm{Gyr}$, at 100 additional outputs with a cadence of $\Delta_{t}=10\,\mathrm{Myr}$ until $t=1\,\mathrm{Gyr}$, and at 99 additional outputs with a cadence of $\Delta_{t}=1\,\mathrm{Gyr}$ until 100 $\mathrm{Gyr}$.
3 Evolution
Figure 1: Evolution of the stellar density $\rho_{*}$ and the velocity anisotropy $\beta_{\mathrm{ani}}$ in simulations, as a function of radius $r$. Left column: One realization of the fiducial system for selected times. Right column: Seven simulations, where $M_{\mathrm{ext}}$ is varied from $M_{\mathrm{IC}}$, at $t=50\,\mathrm{Gyr}$. Top row: Binned $\rho_{*}$ from simulation data. The zoomed-in region shows $\rho_{*}$ for the inner region within $r=0.075\,\mathrm{kpc}$. The dotted line shows $r_{\mathrm{S,P}}=0.2\,\mathrm{kpc}$, and the dashed line shows the analytic Plummer profile $\rho_{\mathrm{P}}$. Middle row: The ratio of $\rho_{*}$ to $\rho_{\mathrm{P}}$, with the dashed line showing $\rho_{*}/\rho_{\mathrm{P}}=1$. Bottom row: Binned $\beta_{\mathrm{ani}}$ from simulation data, with the dashed line showing $\beta_{\mathrm{ani}}=0$. The triangles show the time-average of $\beta_{\mathrm{ani}}$ within (on the left) and outside (on the right) $r_{\mathrm{S,P}}$ for each mass ratio.
Figure 2: Similar to the middle-left panel of Figure 1, but for 10 realizations of the fiducial system, at times $t=0\,\mathrm{Gyr}$ (left) and $t=50\,\mathrm{Gyr}$ (right). The orange line shows the average ratio of all 10 runs.
@@ 3 Evolution
To calculate the evolution of the stellar density profiles from simulation data, we first center the system using the shrinking-sphere method (Power-2003) implemented in pynbody (pynbody) and remove the outermost 250 (5%) of stellar particles. In the outskirts of the system, the sampling is poor, and the estimated density can fluctuate significantly, as the random motions of individual particles can cause a net increase of the maximum radii, and hence, an underestimate of the density. We note that in observed UFDs, the outermost 1–10% of stars are also typically not detectable (e.g., richstein2024) due to the low surface brightness and foreground contamination. We bin the remaining particles in 50 radial shells of equal particle number, each containing 95 particles (95 $\mathrm{M}_{\sun}$). We explicitly test the effect of using different numbers of bins in Appendix A.
@@
Figure 1 shows the evolution of the density and velocity anisotropy profiles of the Plummer stellar system in the external, cuspy gravitational potential. The left column shows the profiles measured from the ICs at $t=0\,\mathrm{Gyr}$ and at seven later times up to $t=100\,\mathrm{Gyr}$, for one fiducial system, evolved in the same external potential that was used to generate the ICs. While the density measured in the simulation fluctuates around the Plummer profile from which the ICs were constructed, there is no discernible evolution. In particular, there is no evidence for a change in the inner density profile induced by the evolution in the cuspy gravitational potential.
@@
The right column shows the profiles measured at $t=50\,\mathrm{Gyr}$, in seven simulations which use identical ICs to the one shown on the left but different external potentials. In this case, the density profile can change, with an increased inner stellar density obtained when the external potential is deeper and a lower one when the external potential is shallower than the potential used to generate the ICs. However, provided that the ICs are created in equilibrium with the external potential, the stellar core is maintained.
@@
As noted by almeida_EIM_2024, circular orbits $(\beta_{\mathrm{ani}}\rightarrow-\infty)$ can accommodate any combination of density and potential, including a cored stellar density in an NFW potential, but those are not physically motivated. Our systems are initially set up with $\beta_{\mathrm{ani}}=0$, and their evolution from this can be seen in the bottom row of Figure 1. As we will show in more detail below, changing the external potential ($M_{\mathrm{ext}}/M_{\mathrm{IC}}\neq 1$) induces net radial motions, which can induce (positive) velocity anisotropy. However, provided that they are evolved in the same potential in which they were created ($M_{\mathrm{ext}}/M_{\mathrm{IC}}=1$), our systems maintain isotropy throughout their evolution.
@@
The fluctuations in the density profile shown in Figure 1 arise simply from the fact that the finite number of $N=5000$ stars provide only a sample of the underlying distribution: observations of UFDs, which often contain only of the order of hundreds of stars (e.g., simon2019), are even more susceptible to sampling variance. This effect is demonstrated in Figure 2, which shows the stellar density profiles of 10 different realizations of the fiducial system, both in the ICs and at $t=50\,\mathrm{Gyr}$. Each individual profile fluctuates around the Plummer value, with $\rho_{*}/\rho_{P}\sim 10\%$. The average of all 10 realizations is close to the Plummer value. There is no discernible evolution, neither in the average nor in any of the individual profiles.
@@
4 Stellar density inference
Figure 4: Results of the stellar density fitting. Top-left panel: The distribution of $\gamma$ for selected times (same as in the left column of Figure 1) for the hierarchical fit. The orange lines show the median of the fit at each time, and the dotted line shows $\gamma=0$. Top-right panel: The PDF of $\gamma$ at $t=0\,\mathrm{Gyr}$. The orange line shows the hierarchical fit, and the remaining lines the non-hierarchical fits for the 10 fiducial simulations. Bottom panels: Correlation of $\gamma$ with the parameters $\alpha$, $\beta$, $r_{\mathrm{S}}$, and $\rho_{\mathrm{S}}$, respectively. The contours show their 68% and 95% confidence intervals, and the dashed crossing lines show the true parameters of the underlying Plummer distribution.

### v2

@@ Abstract
The existence of cuspy or cored centers of dark matter halos is a crucial discriminant between different dark matter models. It has recently been claimed based on dynamical arguments that perfectly cored stellar systems cannot survive inside cuspy dark matter halos, which would make the observation of stellar cores in ultrafaint dwarf galaxies, where dark matter cores cannot form through baryonic processes, a direct falsification of the cold dark matter paradigm. Here, we use idealized simulations to show explicitly that cored stellar systems like those observed in dwarf galaxies can be stable within cuspy dark matter halos over at least several Hubble times. We also demonstrate that observations of ultrafaint dwarf galaxies cannot distinguish mildly positive, flat, or negative inner density slopes, further precluding the dynamical inference of the gravitational potential from the stellar configuration.
@@
software: gadget-4 (Springel et al., 2021), pynbody (Pontzen et al., 2013), NumPy (Harris et al., 2020), SciPy (Virtanen et al., 2020), Matplotlib (Hunter, 2007), Stan (Stan Development Team, 2018), CmdStanPy (Stan Development Team, 2018), Arviz (Kumar et al., 2019).
@@
I Introduction
Dwarf galaxies, and in particular, the dwarf-spheroidal (dSph) and ultrafaint dwarf (UFD) galaxies surrounding the Milky Way, are among the Universe’s most dark matter (DM)-dominated objects. Both in their abundance and in their internal structure, they are key to understanding the nature of DM.
In the standard, cold dark matter (CDM) model, collisionless DM forms so-called “cuspy” halos, characterized by a divergent inner density slope that is well described by a Navarro–Frenk–White (NFW) density profile (Navarro et al., 1997). Observational tests of this prediction on scales of individual galaxies have yielded mixed results. While rotation curves derived from disk galaxies tend to favor “cored” halos of near-uniform density in their central regions (e.g., Oh et al., 2011; Li et al., 2020), the diversity of observed rotation curves includes individual cases that favor NFW-like cusps (e.g., Oman et al., 2015).
@@ I Introduction
Various techniques have been used to infer the mass distributions within “classical” dSphs (stellar masses in the range $\sim 10^{4}\mbox{--}10^{7}\,\mathrm{M}_{\sun}$) from photometric, spectroscopic, and in some cases proper motion data—again with mixed results. Several studies conclude that the most luminous dSphs (e.g., Fornax and Sculptor) harbor large DM cores, based on Jeans models (Read et al., 2019; Hayashi et al., 2020), orbit-based Schwarzschild models (Jardel and Gebhardt, 2012), direct modeling of the phase-space distribution function (DF; Pascale et al., 2018), and statistical analysis of chemodynamically distinct subpopulations (Battaglia et al., 2008; Walker and Peñarrubia, 2011; Amorisco and Evans, 2011; Agnello and Evans, 2012). Others, analyzing the same spectroscopic data sets, claim insufficient constraining power to distinguish DM cores from CDM cusps (e.g., Strigari et al., 2017, 2018; Zhu et al., 2016). Including Hubble Space Telescope proper motion data that resolve the internal velocity dispersions of the Draco dSph, Vitral et al. (2025) constrain any DM core to have radius $\lesssim 1\,\mathrm{kpc}$. Several studies applying homogeneous analyses to all available dSphs conclude that the most luminous ones tend to favor cores, while the least luminous ones favor cusps (e.g., Read et al., 2019; Hayashi et al., 2020, 2023).
@@
This scenario is consistent with results of some cosmological and hydrodynamical simulations, which show that supernova feedback from sufficiently vigorous star formation can alter the structure of CDM halos, inducing the formation of cores in what would originally have been cuspy halos (Navarro et al., 1996; Mashchenko et al., 2008; Pontzen and Governato, 2012) hosting the most luminous dSphs. This process of astrophysical core formation within CDM halos is expected to become inefficient on scales of the smallest and least luminous UFD galaxies, where feedback from the formation of their paltry stellar populations would have little effect on the structure of the host DM halo (e.g., Madau et al., 2014; Di Cintio et al., 2014; Oñorbe et al., 2015; Orkney et al., 2021).
@@
Thus, the detection of cores in UFDs would directly rule out collisionless CDM. However, the fact that UFDs contain so few stars, usually in a single population, makes the analysis that is already difficult in their brighter counterparts even more so (see, e.g., Simon, 2019, for a review).
@@
In a series of papers, Sánchez Almeida et al. (2024a) proposed a new approach to determine the DM density profile of UFDs, requiring only their (projected) surface brightness profiles and applying the Eddington inversion method (Sánchez Almeida et al., 2023) to infer the gravitational potential. They concluded that an isotropic stellar population cannot have a cored density profile within a cuspy DM halo, because it implies that the phase-space DF, derived using the Eddington inversion method, has negative values. In particular, Sánchez Almeida et al. (2024b) claim that the combined surface density profiles of six UFDs require a cored density profile and thus, rule out a cuspy halo, and Sánchez Almeida (2025) explore these in the context of self-interacting DM.
@@
Here, we test this proposition in two parts. First, we explicitly test the stability of cored stellar systems inside cuspy DM halos, using idealized numerical simulations. We find that they can be stable in cuspy halos, provided that the two are initially set up in equilibrium. Second, we test the ability to constrain the stellar density profile of UFDs using a finite number of stars. We find that, in particular, the inner slope is degenerate, even before considering observational limitations.
II Numerical methods
@@
Traditional cosmological simulations lack both the mass and the spatial resolution to simultaneously model the DM halos and resolve the stellar systems of UFDs. Furthermore, they require gravitational softening, which could artificially induce or maintain cores in the stellar components (e.g., Power et al., 2003).
@@
II.1 Initial conditions
We generate the initial conditions (ICs) used to model the dwarf galaxies using the DF method following Hilz et al. (2012). Our fiducial system consists of a stellar component of mass $M_{\star}=5\times 10^{3}\,\mathrm{M}_{\sun}$ embedded in a DM halo with a virial mass of $M_{\mathrm{vir}}=5\times 10^{8}\,\mathrm{M}_{\sun}$. The DM component is modeled by an NFW density profile:
@@ II.1 Initial conditions
$\rho_{\mathrm{S,NFW}}=\frac{M_{\mathrm{vir}}}{4\pi r_{\mathrm{S,NFW}}^{3}}\left[\log(1+c)-\frac{c}{1+c}\right]^{-1},$ (2)
@@
The stellar component is modeled by a Plummer density profile (Binney and Tremaine, 2008):
@@
The DF $f_{i}$ is computed for both components from their respective density profiles $\rho_{i}$ using Eddington’s formula (Binney and Tremaine, 2008):
@@
where $\sigma_{r}$, $\sigma_{\theta}$, and $\sigma_{\phi}$ are the radial, polar, and azimuthal velocity dispersions, respectively, is zero throughout. Although Sánchez Almeida et al. (2024b) argue that stellar cores and cuspy halos are incompatible even when this assumption is relaxed, velocity isotropy is chosen as the most stringent test. For our fiducial system, $\sigma_{r}=\sigma_{\theta}=\sigma_{\phi}\simeq 5.8\,\mathrm{km/s}$, comparable to values observed for UFDs (e.g., Simon, 2019, and references therein).
II.2 Simulations
We run idealized $N$-body simulations using the gadget-4 code (Springel et al., 2021), compiled using the fast multipole method with second-order multipoles, hierarchical time integration, and an external gravitational field created by a static NFW potential. The stellar distribution is set up to be in equilibrium with the NFW halo at $t=0$. In practice, we add to the equations of motion in gadget-4 an additional acceleration term due to an external NFW profile of mass $M_{\mathrm{ext}}=M_{\mathrm{vir}}$:
$\bm{a}_{\mathrm{NFW}}=\frac{GM_{\mathrm{ext}}}{r^{3}}\frac{\frac{r}{r+r_{\mathrm{S,NFW}}}-\log\left(1+\frac{r}{r_{\mathrm{S,NFW}}}\right)}{\log(1+c)-\frac{c}{1+c}}\bm{x},$ (6)
@@ II.2 Simulations
We create and perform 10 different realizations of our fiducial system. The number of realizations is chosen to mimic the number of UFDs used in the analysis of Sánchez Almeida et al. (2024b). To estimate the effect of the external potential, we also vary its virial mass, $M_{\mathrm{ext}}$, (and radius) from that of one of the initial fiducial systems, $M_{\mathrm{IC}}$, and perform seven simulations with an unequal mass ratio $M_{\mathrm{ext}}/M_{\mathrm{IC}}$. In other words, in these simulations, the initial stellar system is explicitly out of equilibrium with the potential in which it will be evolved.
We investigate the effect of softening by performing simulations with softening lengths of 1, 10, and 100 pc. There was no clear distinction between these values, and we chose a softening of 10 pc to be used for this work. The scale at which the gravitational force is affected by softening is $r_{\mathrm{S,G}}=28\,\mathrm{pc}$ (Springel et al., 2021), much less than the size of the stellar core.
Due to the low number density of stars, both the time between stellar collisions and the relaxation time (e.g., Errani et al., 2025) are much longer than the Hubble time, justifying the collisionless approximation implicit in our simulations. We also note that the total kinetic energy of the stars is $\sim 1000$ times below the binding energy of the cusp. Hence, any possible response of the DM to the stars, which our static potential would fail to capture, is guaranteed to be negligible.
To study both the short-time evolution and the long-term stability, the simulations are analyzed at $t=0\,\mathrm{Gyr}$, at 100 additional outputs with a cadence of $\Delta_{t}=10\,\mathrm{Myr}$ until $t=1\,\mathrm{Gyr}$, and at 99 additional outputs with a cadence of $\Delta_{t}=1\,\mathrm{Gyr}$ until 100 $\mathrm{Gyr}$. The dynamical time of stars at the stellar half-mass radius ($\sim 0.27\,\mathrm{kpc}$) in our system is $\sim 25\,\mathrm{Myr}$, much less than the total simulation time.
III Evolution
Figure 1: Evolution of the binned stellar density $\rho_{\star}$, velocity dispersion $\sigma_{r}$, and velocity anisotropy $\beta_{\mathrm{ani}}$ as functions of radius $r$. The dotted line shows $r_{\mathrm{S,P}}=0.2\,\mathrm{kpc}$, and the dash-dotted line shows $r_{\mathrm{S,G}}=0.028\,\mathrm{kpc}$. Left column: one simulation of the fiducial system for selected times. Right column: seven simulations with unequal mass ratios $M_{\mathrm{ext}}/M_{\mathrm{IC}}$ at $t=50\,\mathrm{Gyr}$. First row: stellar density profile, with the zoomed-in region showing the inner region within $r=0.075\,\mathrm{kpc}$, and the dashed line showing the analytic Plummer profile $\rho_{\mathrm{P}}$. Second row: the ratio of $\rho_{\star}$ to $\rho_{\mathrm{P}}$, with the dashed line showing $\rho_{\star}/\rho_{\mathrm{P}}=1$. Third row: velocity dispersion profile, with the triangles showing the time-average of $\sigma_{r}$ within (on the left) and outside (on the right) $r_{\mathrm{S,P}}$. Fourth row: velocity anisotropy profile, with the dashed line showing $\beta_{\mathrm{ani}}=0$, and the triangles showing the time-averaged $\beta_{\mathrm{ani}}$.
Figure 2: Similar to the left panel of the second row of Figure 1, but for 10 realizations of the fiducial system, at times $t=0\,\mathrm{Gyr}$ (left) and $t=50\,\mathrm{Gyr}$ (right). The orange line shows the average ratio of all runs.
@@ III Evolution
To calculate the evolution of the stellar density profiles from simulation data, we first center the system using the shrinking-sphere method (Power et al., 2003) implemented in pynbody (Pontzen et al., 2013) and remove the outermost 250 (5%) of stellar particles. In the outskirts of the system, the sampling is poor, and the estimated density can fluctuate significantly, as the random motions of individual particles can cause a net increase of the maximum radii and hence, an underestimate of the density. We note that in observed UFDs, the outermost 1%–10% of stars are also typically not detectable (e.g., Richstein et al., 2024) due to the low surface brightness and foreground contamination. We bin the remaining particles in 50 radial shells of equal particle number, each containing 95 particles (95 $\mathrm{M}_{\sun}$). We explicitly test the effect of using different numbers of bins in Appendix A.
@@
Figure 1 shows the evolution of the density, velocity dispersion, and velocity anisotropy profiles of the Plummer stellar system in the external, cuspy gravitational potential. The left column shows the profiles measured from the ICs at $t=0\,\mathrm{Gyr}$ and at seven later times up to $t=100\,\mathrm{Gyr}$, for one fiducial system, evolved in the same external potential that was used to generate the ICs. While the density measured in the simulation fluctuates around the Plummer profile from which the ICs were constructed, as seen in the first two rows of the figure, there is no discernible evolution. In particular, there is no evidence for a change in the inner density profile induced by the evolution in the cuspy gravitational potential.
@@
The right column shows the profiles measured at $t=50\,\mathrm{Gyr}$, in seven simulations which use identical ICs to the one shown on the left but different external potentials. In this case, the density profile can change, with an increased inner stellar density obtained when the external potential is deeper and a decreased one when it is shallower than the potential used to generate the ICs. However, provided that the ICs are created in equilibrium with the external potential, the stellar core is maintained.
The left panel in the third row of Figure 1 shows how the velocity dispersion of the fiducial system fluctuates around the mean value of $\sigma_{r}\simeq 5.8\,\mathrm{km/s}$, with a slight decrease towards the outer region of the system, similar to observations of UFDs. As expected, the right panel shows how deeper or shallower external potentials result in increased or decreased stellar velocity dispersions, respectively.
@@
As noted by Sánchez Almeida et al. (2024a), circular orbits $(\beta_{\mathrm{ani}}\rightarrow-\infty)$ can accommodate any combination of density and potential, including a cored stellar density in an NFW potential, but those are not physically motivated. Our systems are initially set up with $\beta_{\mathrm{ani}}=0$, and their evolution from this can be seen in the bottom row of Figure 1. As we will show in more detail below, changing the external potential ($M_{\mathrm{ext}}/M_{\mathrm{IC}}\neq 1$) induces net radial motions, which can induce (positive) velocity anisotropy. However, provided that they are evolved in the same potential in which they were created ($M_{\mathrm{ext}}/M_{\mathrm{IC}}=1$), our systems maintain isotropy throughout their evolution.
@@
The fluctuations in the density profile shown in Figure 1 arise simply from the fact that the finite number of $N=5000$ stars provide only a sample of the underlying distribution: observations of UFDs, which often contain only of the order of hundreds of stars (e.g., Simon, 2019), are even more susceptible to sampling variance. This effect is demonstrated in Figure 2, which shows the stellar density profiles of 10 different realizations of the fiducial system, both in the ICs and at $t=50\,\mathrm{Gyr}$. Each individual profile fluctuates around the Plummer value, with $\rho_{\star}/\rho_{P}\sim 10\%$. The average of all 10 realizations is close to the Plummer value. There is no discernible evolution, neither in the average nor in any of the individual profiles.
@@
IV Stellar density inference
Figure 4: Results of the stellar density fitting. Top-left panel: the distribution of $\gamma$ for selected times (as in the left column of Figure 1) for the hierarchical fit. The orange lines show the median of the fit at each time, and the dotted line shows $\gamma=0$. Top-right panel: the PDF of $\gamma$ at $t=0\,\mathrm{Gyr}$. The orange line shows the hierarchical fit, and the remaining lines show the nonhierarchical fits for the 10 fiducial simulations. Bottom panels: correlation of $\gamma$ with the parameters $\alpha,\,\beta,\,r_{\mathrm{S}},\,\mathrm{and}\,\rho_{\mathrm{S}}$, respectively. The contours show the 68% and 95% confidence intervals, and the dashed crossing lines show the true parameters of the underlying Plummer distribution.